# Re-boosting Self-Collaboration Parallel Prompt GAN for Unsupervised Image Restoration

### v1

@@ 1 Introduction
Figure 1: Comparison of the proposed self-collaboration strategy with the conventional self-ensemble strategy in training complexity, testing complexity, and performance improvement, respectively.
To address the above-mentioned issues, we introduce an innovative unsupervised restoration framework called Re-boosting Self Collaboration Parallel Prompt GAN (RSCP2GAN). The core self-collaboration (SC) strategy provides the framework with an effective self-boosting capability, enabling the restorer obtained from the conventional GAN framework to evolve continuously and significantly. Specifically, it consists of a prompt learning (PL) module and a restorer ($Res$). The SC strategy operates iteratively by replacing the previous, less powerful fixed restorer $\overline{Res}$ in the PL module with the current, more capable $Res$. The updated PL module then generates higher-quality pseudo-degraded images, further enhancing the $Res$ in subsequent iterations. The comparative analysis with the conventional self-ensemble (SE) strategy, shown in Fig. 1, reveals that the SE requires a k-fold testing cost and obtains only marginal improvement. In contrast, our SC can significantly improve the restorer’s ($Res$) performance by over 1.5 dB without incurring extra testing expenses.
Both SE and SC strategies aim to improve $Res$’s performance but from different perspectives: SE focuses on data augmentation during inference, while SC enhances cooperation between the restorer and the generator during training. Building on this, we propose a re-boosting module for SC, termed the Reb-SC strategy, to further enhance the performance of the PL module and $Res$ without increasing model parameters. Extensive experimental results on deraining and denoising tasks confirm the superiority of our method.
@@
    We propose a Self-Collaboration (SC) strategy that significantly enhances the performance of GAN-based restoration frameworks without increasing inference computational complexity. This strategy relies on two key components: the prompt learning (PL) module and the restorer ($Res$). The PL module is the core part of the prompt-guided degraded image generator capable of synthesizing high-quality degraded images.
@@
    We present a re-boosting module to further increase the effectiveness of the proposed SC strategy, i.e., Reb-SC. Similar to the SC strategy, the inference time also does not increase, and the performance of the restorer ($Res$) can be enhanced further.
@@
This work significantly extends our prior work [23]. The key advancements in this paper include: 1) Extension to General Restoration Framework: Our early work focuses only on image denoising tasks. This paper extends the previous denoising model and noise extract (NE) module to the restoration framework and degradation prompt learning (PL) module, resulting in the generalized Self-Collaboration Parallel Prompt GAN (SCP2GAN) framework. 2) Introduction of Reb-SC Strategy: We introduce the Reb-SC strategy to further improve the performance of the PL module without additional parameters and further improve the performance of the restorer. The framework in this work is RSCP2GAN. 3) Extended Experiments and Analysis: We perform extensive experiments on multiple datasets, including deraining (Rain100L [24], Rain12 [25], RealRainL [26]) and denoising (SIDD [27], DND [28], PolyU [29]), and compare our method against state-of-the-art techniques, demonstrating RSCP2GAN’s superior performance.
@@ 2.1 Supervised Image Restoration
The RIDNet method [43] combines synthetic and real images during training to enhance the model generality for denoising. On the other hand, Cheng et al. [44] generate a set of image basis vectors from the noisy input images and reconstruct them from the subspace formed by these basis vectors to obtain image-denoising results. Numerous approaches simultaneously address Gaussian and real-world noise [7, 8]. NAFNet [5] incorporates a series of straightforward but highly effective enhancements, refining the network and fully realizing its performance potential. Recently, a transformer-based framework [6] has been developed, leveraging the advantages of the self-attention strategy while reducing computational complexity.
Image deraining.
Li et al. [26] use real-world rainy video clips to establish a high-quality dataset named RealRainL, consisting of 1,120 high-resolution paired clean and rainy images with low- and high-density rain streaks. On the other hand, SGINet [33] uses high-level semantic information to improve rain removal and RadNet [45] simultaneously removes rain streaks and raindrops. Recently, MIRNet [46] presents an effective feature extraction module to facilitate image restoration and enhancement, and RCDNet [12] proposes an unfolding technique, employing multi-stage training with M-net and B-net to achieve better deraining results. Most recently, NAFNet [5] shows a simple yet effective and efficient approach that achieves state-of-the-art performance for image restoration with using nonlinear activation functions. Restormer [6] utilizes the attention mechanism for deraining and leverages the transformer framework for deraining and optimizing their benefits while reducing computational complexity. A multi-experts-based DRSFormer [35] provides more accurate detail and texture recovery. However, the number of real-world degraded images is limited, and creating large amounts of paired training data is time-consuming and labor-intensive.
@@ 2.2 Unsupervised Image Restoration
Numerous restoration methods using a large amount of synthetic data from generative adversarial networks for training have been developed.
@@
GCBD [16] uses a generator capable of producing pseudo-noisy images to train a denoiser, and CycleGAN [47] is introduced for further improvement. Among these methods, GAN2GAN [22] uses a multi-generator / discriminator architecture to enhance the extraction of noisy information and generate synthetic images that closely match the real noise distribution. On the other hand, Hong et al. [17] introduce UIDNet, which utilizes a sharpening processing mechanism to achieve noise separation and improve the training of unpaired denoising models. In SCPGabNet [23], a self-collaboration strategy that iteratively enhances the performance of the denoising network has been proposed, leading to significant improvements over conventional GAN frameworks. Additionally, various methods have emerged for training models exclusively with noisy images, called self-supervised denoising. CVF-SID [48] integrates cyclic adversarial learning with the self-supervised residual framework. Recently, a self-supervised framework named AP-BSN [49] has been shown to effectively manage real-world signal-dependent noise and adapt well to realistic noise conditions. Most recently, LG-BPN [50] shows masking the central region of a large convolution kernel to reduce the spatial correlation of noise and introduces a dilated Transformer block to capture global information, while others are developed to APBSN using random sampling for augmentation [51].
Image deraining.
Numerous unsupervised image restoration methods have been developed [52, 53, 54, 20] based on the CycleGAN model. The DerainCycleGAN [20] extracts the rain streak masks using two constrained cycle-consistency branches by paying attention to both the rainy and rain-free image domains for restoration. Yu et al. [55] consider the prior knowledge of the rain streak and connect the model-driven and data-driven methods via an unsupervised learning framework. DCD-GAN [21] incorporates contrastive learning loss as a constraint during network training, which enhances the model performance. NLCL [56] uses a decomposition-based non-local contrastive learning strategy to compute the self-similarity of the image for restoration.
@@ 3.1.1 Prompt-Guided Degraded Image Generator
Despite the state-of-the-art performance of supervised image restoration frameworks on numerous benchmark datasets, they require a large amount of paired data, which is challenging to obtain in real-world scenarios. Although several GAN-based unsupervised frameworks can address this limitation, the performance is often inferior to supervised approaches. This performance gap is largely due to the domain gap between synthetic and real-world degraded images [22, 57]. To bridge this gap, the key is to enhance the quality of synthetic images to make them as close as possible to real-world degraded images. If synthetic images can be generated to closely resemble real degraded images, the performance of unsupervised restoration models can approach the effectiveness of their supervised counterparts. Improving the quality of synthetic images is thus critical to boosting the performance of unsupervised restoration frameworks.
@@ 3.1.3 GAN-based Degradation Synthesis
which means for the generated image $x_{u-syn}$, its adversarial loss $L_{\text{adv2}}$ is constrained between $y$ and $x_{u-syn}$. The other three adversarial losses can be constructed similarly by constraining the current generated images and $y$. The overall loss for the GAN model is:
@@ 3.2.1 Self-Collaboration Strategy
Figure 6: Illustration of the Reb-SC strategy. The self-ensemble (SE) can improve the model’s performance during testing but increases the network’s computational complexity. In contrast, our self-collaboration (SC) is used only in the training phase, does not increase inference time, and can significantly enhance the restorer’s performance. Based on this, we combine the characteristics of the SE and SC to propose Reb-SC. In Stage 1, we use the original SC; in Stage 2 and Stage 3, we use Reb-SC. We input augmentation of low-quality images into the fixed PL module to enhance the performance of the $\overline{Res}$ in PL, thereby training a better restorer $Res$. The SE can bring a 0.12 dB improvement to the $Res$, while our SC can provide a 1.61 dB improvement. However, the Reb-SC combines both characteristics and can achieve a 1.93 dB improvement.
@@ 3.2.2 Re-boosting SC Strategy
Figure 7: (a) Visualization of the unpaired synthetic degraded image by DCD-GAN [21] and our RSCP2GAN. (b) Quantitative comparison between DCD-GAN and our framework for generator and restorer.
@@ 3.2.3 Analysis of RSCP2GAN
TABLE I: Denoising results of several competitive methods on SIDD Validation, SIDD Benchmark, and DND Benchmark. The highest PSNR(dB)/SSIM is highlighted in red, while the second is in blue. Additionally, * denotes that the approach is trained on the DND benchmark directly, and the results without * means the methods trained on the SIDD datasets.
@@
TABLE II: Quanlitative Comparison on the PolyU Dataset.

### v2

@@ 1 Introduction
Figure 1: Comparison of the proposed self-collaboration (SC) and re-boosting SC (Reb-SC) strategies with the conventional self-ensemble (SE) strategy in training expenses, testing expenses, and performance improvement, respectively.
To address the above-mentioned issues, we introduce an innovative unsupervised restoration framework called Re-boosting Self Collaboration Parallel Prompt GAN (RSCP2GAN). The core self-collaboration (SC) strategy provides the framework with an effective self-boosting capability, enabling the restorer obtained from the conventional GAN framework to evolve continuously and significantly. Specifically, it consists of a prompt learning (PL) module and a restorer ($Res$). The SC strategy operates iteratively by replacing the previous, less powerful fixed restorer $\overline{Res}$ in the PL module with the current, more capable $Res$. The updated PL module then generates higher-quality pseudo-degraded images, further enhancing the $Res$ in subsequent iterations. The comparative analysis with the conventional self-ensemble (SE) strategy is shown in Fig. 1. Traditional SE is applied during the testing phase of a trained model, where the input is augmented, followed by multi-branch restoration and averaging. This significantly increases both the testing time and memory usage, which is disadvantageous when applying restoration networks in the real world. Meanwhile, it obtains only a marginal improvement. In contrast, our SC can significantly improve the restorer’s ($Res$) performance by over 1.5 dB without incurring extra testing expenses.
Both SE and SC strategies aim to improve $Res$’s performance but from different perspectives: SE focuses on data augmentation during inference, while SC enhances cooperation between the restorer and the generator during training. Building on this, we propose a re-boosting module for SC, termed the Reb-SC strategy. The re-boosting applies data augmentation during the training phase in our SC strategy, further enhancing the proposed SC. The Reb-SC, like the original SC, is applied during the training phase, improving the performance of the restorer significantly and making only minor changes to the framework, without affecting testing time or memory usage, and having no impact on the application of restoration networks. By this way, our SC can further improve the performance of the restorer. Please see Sec. 3.2.4 for more detailed analysis. Extensive experimental results on denoising, deraining, and desnowing tasks confirm the superiority of our method.
@@
    We propose a Self-Collaboration (SC) strategy that significantly enhances the performance of GAN-based restoration frameworks without increasing inference computational complexity. This strategy relies on two key components: the prompt learning (PL) module and the restorer ($Res$). The PL module is the core part of the prompt-guided degraded image generator, capable of synthesizing high-quality degraded images.
@@
    We present a re-boosting module to increase further the effectiveness of the proposed SC strategy, i.e., Reb-SC. Similar to the SC strategy, the inference time also does not increase, and the performance of the restorer ($Res$) can be enhanced further.
@@
This work significantly extends our prior work [23]. The key advancements in this paper include: 1) Extension to General Restoration Framework: Our early work focuses only on image denoising tasks. This paper extends the previous denoising model and noise extract (NE) module to the restoration framework and degradation prompt learning (PL) module, resulting in the generalized Self-Collaboration Parallel Prompt GAN (SCP2GAN) framework. 2) Introduction of Reb-SC Strategy: We introduce the Reb-SC strategy to further improve the performance of the PL module without additional parameters and further improve the performance of the restorer. The framework in this work is RSCP2GAN. 3) Extended Experiments and Analysis: We perform extensive experiments on multiple datasets, including deraining (Rain100L [24], Rain12 [25], RealRainL [26]), desnowing (CSD [27] and Snow100K [28]) and denoising (SIDD [29], DND [30], PolyU [31]), and our method performs well than state-of-the-art techniques, demonstrating RSCP2GAN’s superior performance. 4) Analysis of Training/Testing Expenses and Strategy Effectiveness: We add more analyses about training expenses, testing expenses, and performance improvement on the proposed SC&Reb-SC strategies and traditional self-ensemble augmentation to clarify our contribution.
@@ 2.1 Supervised Image Restoration
The RIDNet method [45] combines synthetic and real images during training to enhance the model’s generality for denoising. On the other hand, Cheng et al. [46] generate a set of image basis vectors from the noisy input images and reconstruct them from the subspace formed by these basis vectors to obtain image-denoising results. Numerous approaches simultaneously address Gaussian and real-world noise [7, 8]. NAFNet [5] incorporates a series of straightforward but highly effective enhancements, refining the network and fully realizing its performance potential. Recently, a transformer-based framework [6] has been developed, leveraging the advantages of the self-attention strategy while reducing computational complexity.
Image deraining and desnowing.
As common forms of weather-induced degradation, rain and snow have received considerable attention in image restoration. To tackle this challenge, numerous studies have introduced learning-based approaches that demonstrate strong performance in restoring images degraded by rain or snow. For image deraining, Li et al. [26] use real-world rainy video clips to establish a high-quality dataset named RealRainL, consisting of 1,120 high-resolution paired clean and rainy images with low- and high-density rain streaks. On the other hand, SGINet [35] uses high-level semantic information to improve rain removal and RadNet [47] simultaneously removes rain streaks and raindrops. Recently, MIRNet [48] presents an effective feature extraction module to facilitate image restoration and enhancement, and RCDNet [12] proposes an unfolding technique, employing multi-stage training with M-net and B-net to achieve better deraining results. Most recently, NAFNet [5] shows a simple yet effective and efficient approach that achieves state-of-the-art performance for image restoration with using nonlinear activation functions. Restormer [6] utilizes the attention mechanism for deraining and leverages the transformer framework for deraining and optimizing their benefits while reducing computational complexity. A multi-expert-based DRSFormer [37] provides more accurate detail and texture recovery. For image desnowing, Zhang et al. [49] present a densely connected multi-scale architecture that leverages semantic segmentation and depth priors via a self-attention mechanism to enhance image restoration performance. Chen et al. [27] propose a single-image desnowing framework that integrates a hierarchical dual-tree complex wavelet transform for better representation of snow particles of various shapes and scales. They further introduced a novel, contrast channel loss to exploit the intensity differences between snowy and clean regions, significantly improving snow removal performance. Quan et al. [50] propose InvDSNet, a dual-path invertible neural network for single image desnowing, which enables effective snowflake removal while preserving image details through progressive feature disentanglement and invertible reconstruction. Meanwhile, PromptIR [51] presents a prompt-learning-based framework for all-in-one blind image restoration, aiming to handle various degradation types, including deraining, within a unified model. The method introduces a lightweight Prompt Block module that generates input-conditioned prompts to encode degradation-specific cues, which are then injected into multiple stages of the transformer decoder to guide the restoration process. After that, LMQFormer [52], a lightweight yet effective snow removal network, is proposed, which leverages a Laplace-guided VQVAE to generate a coarse snow mask and a Mask Query Transformer to refine the restoration. By introducing duplicated mask query attention, the model focuses computation on snow regions, achieving state-of-the-art performance with minimal parameters and fast inference speed. To better model the spatially-varying and multi-scale nature of rain streaks, Chen et al. [53] develop a bidirectional multi-scale Transformer architecture augmented with implicit neural representations (INRs). Unlike previous single-scale Transformer-based methods that process rain effects at a fixed resolution, their model leverages both coarse-to-fine and fine-to-coarse feature propagation to exploit cross-scale dependencies. More recently, PEUNet [54] incorporates physical priors—such as atmospheric light and snow shape—to refine the desnowing process iteratively. By formulating snow removal as an optimization problem under a MAP framework and introducing a novel snow shape prior as surrogate supervision, their method enhances both interpretability and performance, achieving state-of-the-art results across multiple benchmarks.
However, the number of real-world degraded images is limited, and creating large amounts of paired training data is time-consuming and labor-intensive.
@@ 2.2 Unsupervised Image Restoration
In unsupervised settings, due to the lack of real paired datasets, synthetic images are commonly used to approximate real training data. Thus, it is crucial to address the domain gap between the synthetic and real image domains. Fu et al. [55] introduce a CNN-based reconstruction method leveraging both external priors and image-specific internal learning, demonstrating strong performance and generalization on real-world coded hyperspectral data. Da et al. [56] propose PromptGAT, a novel sim-to-real transfer framework that leverages prompt-based learning with large language models to enhance the adaptability of reinforcement learning policies to the real-world traffic signal control task. Meanwhile, numerous unsupervised restoration methods have been developed using a large amount of synthetic data from generative adversarial networks to train the model. In the following sections, we review how prior works tackle the domain gap and enhance the performance of restoration networks under unsupervised settings in image denoising, deraining, and desnowing tasks.
@@
GCBD [16] uses a generator capable of producing pseudo-noisy images to train a denoiser, and CycleGAN [57] is introduced for further improvement. Among these methods, GAN2GAN [22] uses a multi-generator / discriminator architecture to enhance the extraction of noisy information and generate synthetic images that closely match the real noise distribution. On the other hand, Hong et al. [17] introduce UIDNet, which utilizes a sharpening processing mechanism to achieve noise separation and improve the training of unpaired denoising models. In SCPGabNet [23], a self-collaboration strategy that iteratively enhances the performance of the denoising network has been proposed, leading to significant improvements over conventional GAN frameworks. Additionally, various methods have emerged for training models exclusively with noisy images, called self-supervised denoising. CVF-SID [58] integrates cyclic adversarial learning with the self-supervised residual framework. Recently, a self-supervised framework named AP-BSN [59] has been shown to effectively manage real-world signal-dependent noise and adapt well to realistic noise conditions. Most recently, LG-BPN [60] shows masking the central region of a large convolution kernel to reduce the spatial correlation of noise and introduces a dilated Transformer block to capture global information, while others are developed for APBSN using random sampling for augmentation [61].
Image deraining and desnowing.
Numerous unsupervised image restoration methods have been developed [62, 63, 64, 20] based on the CycleGAN model. The DerainCycleGAN [20] extracts the rain streak masks using two constrained cycle-consistency branches by paying attention to both the rainy and rain-free image domains for restoration. Yu et al. [65] consider the prior knowledge of the rain streak and connect the model-driven and data-driven methods via an unsupervised learning framework. Yasarla et al. [66] introduce a weather-agnostic unsupervised restoration approach by augmenting CycleGAN with Deep Gaussian Process-based latent supervision, enabling effective training on unpaired data across diverse weather conditions. In addition, Xie et al. [67] propose UAIR, an unsupervised all-in-one adverse weather image restoration framework that leverages contrastive learning to enhance both content preservation and category alignment without requiring paired training data. DCD-GAN [21] incorporates contrastive learning loss as a constraint during network training, which enhances the model performance. NLCL [68] uses a decomposition-based non-local contrastive learning strategy to compute the self-similarity of the image for restoration.
@@ 3.1.1 Prompt-Guided Degraded Image Generator
Despite the state-of-the-art performance of supervised image restoration frameworks on numerous benchmark datasets, they require a large amount of paired data, which is challenging to obtain in real-world scenarios. Although several GAN-based unsupervised frameworks can address this limitation, the performance is often inferior to that of supervised approaches. This performance gap is largely due to the domain gap between synthetic and real-world degraded images [22, 69]. To bridge this gap, the key is to enhance the quality of synthetic images to make them as close as possible to real-world degraded images. If synthetic images can be generated to closely resemble real degraded images, the performance of unsupervised restoration models can approach the effectiveness of their supervised counterparts. Improving the quality of synthetic images is thus critical to boosting the performance of unsupervised restoration frameworks.
@@ 3.1.3 GAN-based Degradation Synthesis
which means for the generated image $x_{u-syn}$, its adversarial loss $L_{\text{adv2}}$ is constrained between $y$ and $x_{u-syn}$. The other three adversarial losses can be constructed similarly by constraining the current generated images and $y$:
$L_{\text{adv1}}$ $=\|D(y)-\mathbf{1}\|_{2}^{2}+\|D(y_{s\text{-}syn})-\mathbf{0}\|_{2}^{2}$ (3)
$L_{\text{adv3}}$ $=\|D(y)-\mathbf{1}\|_{2}^{2}+\|D(y_{u\text{-}syn})-\mathbf{0}\|_{2}^{2}$
$L_{\text{adv4}}$ $=\|D(y)-\mathbf{1}\|_{2}^{2}+\|D(x_{s\text{-}syn})-\mathbf{0}\|_{2}^{2}$
The overall loss for the GAN model is:
@@ 3.2.1 Self-Collaboration Strategy
Figure 6: Illustration of the Reb-SC strategy. The self-ensemble (SE) can improve the model’s performance during testing, but it increases the network’s computational complexity. In contrast, our self-collaboration (SC) is used only in the training phase, does not increase inference time, and can significantly enhance the restorer’s performance. Based on this, we combine the characteristics of the SE and SC to propose Reb-SC. In Stage 1, we use the original SC; in Stage 2 and Stage 3, we use Reb-SC. We input augmentation of low-quality images into the fixed PL module to enhance the performance of the $\overline{Res}$ in PL, thereby training a better restorer $Res$. The SE can bring a 0.12 dB improvement to the $Res$, while our SC can provide a 1.61 dB improvement. However, the Reb-SC combines both characteristics and can further improve the performance of the restorer, achieving a 1.93 dB improvement.
@@ 3.2.2 Re-Boosting SC Strategy
Figure 7: (a) Visualization of the unpaired synthetic degraded image by DCD-GAN [21] and our RSCP2GAN. (b) Quantitative comparison between DCD-GAN and our framework for the generator and the restorer.
@@ 3.2.3 Deep Analysis of RSCP2GAN
TABLE I: Denoising results of several competitive methods on SIDD Validation, SIDD Benchmark, and DND Benchmark. Additionally, * denotes that the approach is trained on the DND benchmark directly, and the results without * means the methods trained on the SIDD datasets.
@@
TABLE II: Qualitative Comparison on the PolyU Dataset.
@@
3.2.4 Analysis of the Proposed SC and Reb-SC
Traditional data augmentation is applied during the testing phase of a trained model, where data is augmented, followed by multi-branch restoration and averaging. This significantly increases both the testing time and memory usage, which is disadvantageous when applying restoration networks in the real world. In contrast, our re-boosting applies data augmentation during the training phase, further enhancing the proposed SC. Like the original SC, it is applied during the training phase, making only minor changes to the framework, without affecting testing time or memory usage in the reference phase, and having no impact on the application of restoration networks.
Training: Taking the original framework without self-collaboration as a reference, suppose it requires training for $T_{0}$ epochs, with a per-epoch complexity of $P_{0}$, which contains the complexity from the generator, discriminator, and restorer. The total complexity of training the original framework $P_{Ori_{Train}}=T_{0}\times P_{0}$.
After applying SC, the training runs for additional $T_{s}$ epochs ($T_{s}\ll T_{0}$), with each epoch having a complexity of $P_{s}$. Since the parameter $\overline{Res}$ in the SC is fixed, and is larger than the initial simple Gaussian filter in the basic stage without SC of the P2GAN (See Algorithm 1), while the complexity of the generator and discriminator is not changed, it only adds about (slightly lower) $\overline{Res}$’s complexity. Therefore, we get $P_{s}$ higher than $P_{0}$, but much less than $2\times P_{0}$, i.e., $P_{0}<P_{s}<2\times P_{0}$. Thus, the complexity in the SC phase is $P_{SC_{Train}}=T_{s}\times P_{s}$, where $T_{s}\times P_{0}<P_{SC_{Train}}<2\times T_{s}\times P_{0}$. As shown in Fig. 6, for the Reb-SC, we have two or four input images (the degraded images with different augmentation) to the $\overline{Res}$, and run only one epoch in each stage. Thus, the complexity $P^{\prime}_{s}$ of the framework in Reb-SC is less than (4 + 2) ×$\times P_{s}$ = 6 ×$\times P_{s}$ and larger than $2\times P_{0}$. This introduces additional complexity for the Reb-SC stage, denoted as $P_{REB_{Train}}$, such that: $2\times P_{0}<P_{REB_{Train}}<6\times P_{s}<2\times 6\times P_{0}=12\times P_{0}$. Thus, the overall training complexity of the framework becomes $P_{Train}=P_{Ori_{Train}}+P_{SC_{Train}}+P_{REB_{Train}}$, where $T_{0}\times P_{0}+T_{s}+2\times P_{0}=(T_{0}+T_{s}+2)\times P_{0}<P_{Train}<T_{0}\times P_{0}+2\times T_{s}\times P_{0}+12\times P_{0}=(T_{0}+2\times T_{s}+12)\times P_{0}$. Since ($122\times T_{s}+$ 12) is relatively small compared to $T_{0}$, the added complexity $(2\times T_{s}+12)\times P_{0}$ is also smaller than the original complexity $T_{0}\times P_{0}$, resulting in a slight increase in training complexity.
Inference: The same restoration module is used as in the non-Reb-SC version, so the inference complexity is simply that of a single denoising pass, denoted as $P_{SC\&REB_{Inference}}=P_{Inference}$. In traditional self-ensemble augmentation strategies, $N$ augmented variants are typically averaged during inference, leading to a complexity of $P_{Aug_{Inference}}=N\times P_{Inference}$ $N\times P_{0}$ complexity. When $N=8$, such traditional self-ensemble augmentation significantly increases the inference complexity to $8\times P_{0}$.
Performance Improvement: As we presented in Fig. 7, under the CycleGAN framework, the performance of the restorer depends on the quality of the data generated by the generator — the better the generated data, the better the trained restorer. There is a strong positive correlation between them. The quality of the generator, in turn, depends on the accuracy of the degradation hint information in the input. Taking denoising as an example, this refers to the accuracy of noise estimation. In our framework, we compute the noise information by subtracting the clean image (restored by the current restorer) from the real noisy image. Thus, the better the current restoration, the more accurate the extracted noise information becomes. More accurate noise information leads to more precise pseudo-noisy images, which in turn further improve the performance of the restorer. Consequently, the denoising performance will gradually improve with our method. The performance improvement analyses of other restoration tasks using our framework are similar.
Specifically, let $Q_{t}$ be the current restoration performance of the restorer in our framework, $Q_{t-1}$ and $Q_{t+1}$ denote the previous and next SC/Reb-SC iteration performance of the restorer. Let $P_{t}$ be the current quality of the input prompt to the generator, which is obtained by the degraded image subtract the previous restored image by the restorer with restoration performance $Q_{t-1}$. Let $G_{t}$ be the current performance of the generator. Firstly, we have $Q_{0}$, $P_{0}$, and $G_{0}$, where $Q_{0}$ is the performance of a simple filter (as illustrated in Sec. 3.2.1 and Algorithm 1). When applying SC, we use a better restorer with $Q_{1}$ ($Q_{1}$ refers to the performance of the restoration network NAFNet trained under the complete unsupervised framework without SC or Reb-SC. Thanks to NAFNet’s advanced design and its deep exploitation of data characteristics, it achieves superior restoration performance.) to replace the previous filter with $Q_{0}$ ($Q_{0}$ represents the performance of a simple Gaussian filter with a basic convolutional structure and limited fitting capacity.) in the framework. Then we get a better prompt $P_{1}$, which can further train a better generator with $G_{1}$. After this SC iteration, we get a better restorer with $Q_{2}$. Following the same process, we then get an improved prompt $P_{2}$ and generator with $G_{2}$. Similarly, we can conclude that if the restorer with $Q_{T_{SC}}$ is better than $Q_{T_{SC}-1}$ ($Q_{T_{SC}-1}$ is used in the input of the generator with $G_{T_{SC}-1}$, and $T_{SC}$ is the current iteration number of the SC phase), replacing the restorer from $Q_{T_{SC}-1}$ to $Q_{T_{SC}}$ can generate better prompt $P_{T_{SC}}$ (i.e., $P_{T_{SC}}>P_{T_{SC}-1}$). A better $P_{T_{SC}}$ leads to a stronger generator with $G_{T_{SC}}$ (i.e., $G_{T_{SC}}>G_{T_{SC}-1}$), thus improving the restoration performance. By continuously iterating in this manner, the performance improves until convergence.
At the Reb-SC stage, we have two or four input images (the degraded images with different augmentations) to the $\overline{Res}$ and average the restored results to get a better restored image. Let $T_{Reb-SC}$ be the current iteration number of the Reb-SC phase. Then, the input of the generator with $G_{T_{Reb-SC}-2}$ will be improved since multiple restored results generated by the restorer with $Q_{T_{Reb-SC}-1}$ are averaged, leading to a better $P_{T_{Reb-SC}-1}$. The better $P_{T_{Reb-SC}-1}$ will enhance the generator further with $G_{T_{Reb-SC}-1}$, and gets a better restorer $Q_{T_{Reb-SC}}$. After this process, we get the final restorer with the best performance.